OpenStudy (anonymous):

HELP A zoo train ride costs $3 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who took the ride was 30, and the total money collected was $50. What was the number of children and the number of adults who took the train ride that day, and which pair of equations can be solved to find the numbers?

OpenStudy (anonymous):

@joeycoulter76 @JoannaBlackwelder @MiningProMC @beezle1197 @ospreytriple

OpenStudy (anonymous):

ANSWER CHOICES 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 3a + c = 50 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 3a − c = 50 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 3a − c = 50 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 3a + c = 50

OpenStudy (texaschic101):

total of adults and children was 30 : a + c = 30 $3 per adult, $1 per child....total of $50 : 3a + c = 50 a + c = 30 a = 30 - c -- now sub 30 - c in for a in the other equation 3a + c = 50 3(30 - c) + c = 50 -- distribute through the parenthesis 90 - 3c + c = 50 -- subtract 90 from both sides -3c + c = 50 - 90 -- simplify -2c = - 40 -- divide both sides by -2 c = 20 a + c = 30 a + 20 = 30 a = 30 - 20 a = 10 so there are 20 children and 10 adults

OpenStudy (anonymous):

so a or c?

OpenStudy (texaschic101):

A

OpenStudy (anonymous):

thank u, can u do another plz

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

No direct answers, please.

OpenStudy (anonymous):

How many solutions are there to the given set of equations? Equation C: y = 3x + 7 Equation D: y = 3x + 2 How many solutions are there to the given set of equations? One solution Two solutions Infinitely many solutions No solution

OpenStudy (texaschic101):

ok...lets do this one together y = 3x + 7......so sub in 3x + 7 for y in the other equation 3x + 7 = 3x + 2 now do you know how to solve for y ??

OpenStudy (anonymous):

im not super sure

OpenStudy (texaschic101):

ooops...I mean for x

OpenStudy (texaschic101):

3x + 7 = 3x + 2 --- subtract 7 from both sides 3x + 7 - 7 = 3x + 2 - 7 -- simplify 3x = 3x - 5 -- now subtract 3x from both sides 3x - 3x = 3x - 3x - 5 -- simplify 0 = -5 now...is that correct ? Does 0 = -5 ??

OpenStudy (anonymous):

yes?

OpenStudy (texaschic101):

no.....0 does not equal -5....therefore, there are no solutions to this problem

OpenStudy (anonymous):

ohh duh i see now, thank u

OpenStudy (texaschic101):

sure thing :)